# Seeing the Forest for the Trees: A Large Scale, Continuously Updating Meta-Analysis of Frontier LLMs

### v1

@@ 4.2 Which Tasks Benefit from ICL?
Previous research has systematically explored ICL behavior (Min et al., 2022; Agarwal et al., 2024; Wei et al., 2023; Bertsch et al., 2024). In line with our study in section 4.1, we analyze the improvement of using in-context examples compared to not using any. To achieve this, we extract instances from the same papers and tables that compare few-shot and zero-shot setups under identical conditions (i.e., model, dataset, subset, metric, and prompting method). In this case, no manual labor was needed beyond implementation, as the number of in-context examples is an integer value that can be easily computed for filtering. We measure performance deltas between few-shot and zero-shot setups, where positive values indicate few-shot advantages and negative values indicate zero-shot advantages.
Figure 7: Few-shot learning significantly improves performance over zero-shot learning in multimodal tasks, whereas its impact on math tasks is less pronounced compared to CoT, as shown in Fig. 5. Grey dots represent individual deltas (improvements), sky pink dots show the mean delta per paper, and a pink star marks the mean delta for each category.
@@
Figure 7 presents the results. In contrast to CoT’s strong performance in mathematical reasoning, ICL shows only modest benefits for math tasks (mean delta 11.1 vs mean delta 2.6). However, ICL demonstrates more substantial improvements in multimodal applications, despite showing considerable performance variability across different cases. This variance may be due to the broad scope of multimodal tasks, including speech and image processing. ICL shows more uniform improvements over zero-shot performance across different categories compared to the study in Fig. 5 (std: 8.3). Note that we did not account for the number of in-context examples used; we only distinguished between the use of in-context examples and zero-shot examples. We also plot the performance improvement distribution of more versus fewer in-context examples in appendix C and find a similar overall median. This suggests that the presence of demonstrations is more important than their quantity.
@@ 4.3 Which Dataset Characteristics Negatively Affect CoT and ICL?
Motivation and Setup
To comprehend the patterns of performance degradation, we analyze cases where performance declined: specifically, where CoT resulted in worse outcomes than direct prompting, and where few-shot learning performed worse than zero-shot learning (i.e. deltas below zero). Table 4 summarizes the key dataset characteristics associated with performance declines under both approaches based on analyzing the descriptions.
Analysis
Out of performance decline cases, tasks requiring expert-level knowledge showed the highest ratio for both cases, approximately 31%, indicating that knowledge-intensive tasks do not benefit much from different prompting configurations alone. Sprague et al. (2024) notes a similar point that apparent performance gains in knowledge tasks mainly stem from "reasoning" or "math" components in datasets like MMLU (Hendrycks et al., 2020). This also aligns partly with Liu et al. (2024), who found that CoT can hinder LLM performance on tasks requiring minimal deliberative thinking.
@@
4.4 How do CoT and ICL Interact?
Setup and Motivation
Building on previous studies section 4.1 and section 4.2, our analysis uncovers how CoT reasoning interacts with demonstrations in prompting, highlighting their complementarity. We extract instances where CoT is used with varying levels of demonstrations, as well as instances that compare CoT and direct prompting under different demonstration levels (zero-shot vs. few-shot), all while keeping other attributes constant.
@@ 4.4 How do CoT and ICL Interact?
From Table 5, demonstrations consistently enhance CoT performance: few-shot CoT outperforms zero-shot CoT by 3.0 points, and increasing the number of demonstrations yields a similar 3.1 points improvement. These comparable gains suggest that the mere presence of demonstrations, rather than their quantity, drives the improvement. However, from Table 6, we can see that when comparing CoT against standard prompting, the benefits remain relatively constant regardless of demonstration use. CoT’s improvement over standard prompting is similar in both few-shot and zero-shot settings. This indicates that while demonstrations enhance both approaches, they do not amplify CoT’s relative advantage over standard prompting.
@@ 5 Related Work
Previous works have focused on extracting basic result tuples (e.g., task, dataset, metric) from scientific literature (Singh et al., 2019; Hou et al., 2019; Kardas et al., 2020; Yang et al., 2022; Bai et al., 2023; Singh et al., 2024; Şahinuç et al., 2024; Kabongo et al., 2024). Our extraction pipeline improves upon this approach in two significant ways: it extracts enriched tuples that include prompting-related attributes and generates detailed dataset descriptions by leveraging LLM and automatically linked source papers. Hence, unlike previous works that primarily compiled leaderboard tables, our enhanced extraction pipeline enables deeper meta-analysis, contributing to the broader goal of AI-driven scientific discovery (Xu et al., 2021; Majumder et al., 2024; M. Bran et al., 2024).
@@
Our study focuses on extracting experimental results of frontier proprietary LLMs (Achiam et al., 2023; Anthropic, 2024; Team et al., 2023), with a specific emphasis on target attributes that incorporate information about prompting methods (Brown et al., 2020; Wei et al., 2022; Min et al., 2022). In the context of prompting, prior studies have systematically analyzed the mechanisms behind prompting methods, focusing either on the use of in-context examples (Min et al., 2022; Lampinen et al., 2022; Weber et al., 2023; Zhang et al., 2022) or techniques that elicit reasoning, such as CoT prompting (Wei et al., 2022; Shaikh et al., 2023; Wang et al., 2023; Turpin et al., 2024; Sprague et al., 2024; Liu et al., 2024). In contrast, we analyze the model’s behavior through meta-analysis which aggregates data from scientific literature, identifying trends and linking resources for deeper investigation.
Meta Analysis
Meta-analysis systematically aggregates and examines data from multiple independent studies on a given topic to derive more precise and reliable conclusions. It has been widely applied in the biomedical domain for identifying target materials or clinical records (Bao et al., 2019; Yun et al., 2024). In the NLP domain, meta-analysis has been used for metric standardization (Reiter, 2018), literature review (Santu et al., 2024; Du et al., 2024), and assessing evaluation criteria for specific domains (Ostheimer et al., 2023). In contrast, our work employs a meta-analysis approach to evaluate the behavior of LLMs. In the context of LLMs, Asai et al. (2024) utilizes retrieval-augmented language models to synthesize scientific literature. However, this approach can only process a limited number of documents during retrieval to synthesize. The work by Sprague et al. (2024) is perhaps the most closely related to ours. They conducted a meta-analysis through a literature survey to examine the effectiveness of CoT prompting for LLMs. However, their study is primarily focused on CoT prompting, conducted on a limited scale, and relies on manual extraction methods.
@@ 6 Conclusion
Our study streamlines meta-analysis by using an LLM for dataset extraction, demonstrating that automatically extracted datasets can yield novel findings and replicate manual analyses. We confirm our dataset’s quality by replicating a key finding from Sprague et al. (2024) on CoT. Our analysis provides insights into prompting configurations, such as the benefits of in-context examples, the combined effects of CoT and in-context examples, and the dataset characteristics on performance declines with CoT or demonstrations. Overall, our resources support ongoing meta-analyses, enhancing the understanding of LLM behaviors and trends.
@@ 7 Limitations
Our study focuses on four leading proprietary LLMs, selected for their widespread adoption and extensive documentation in existing literature. We excluded newer models like GPT4-o1 and Deepseek-R1 due to limited published results, though our analysis pipeline can easily accommodate them as more experimental data becomes available. While most analyzed models are not fine-tunable, this limitation presents an opportunity for future research combining fine-tuning and prompting methods analysis.
@@
Our meta-analysis aims to identify trends by aggregating information across multiple studies, which generates potential hypotheses but requires systematic validation. While we report findings and patterns observed in the aggregated scientific literature, we did not independently validate each claim or finding. This limitation suggests the need for future work to rigorously test the hypotheses emerging from our analysis.
@@ Acknowledgments
We would like to thank Microsoft’s Azure Accelerate Foundation Models Research Program and NVIDIA’s Academic Grant Program for providing computational resources to support this work. This research is supported in part by the NSF under grant number IIS-2052498 and SMA-2418946. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. We also appreciate Ethan Mendes, Duong Minh Le, and Seongeun Park for their valuable discussions and feedback on the paper.
@@ Appendix A Details and Trials During Extraction
Before the extraction stage, we filter tables to reduce extraction time and API costs. To cut computational expenses, we pre-filter tables for those containing results of target models. Using simple heuristics, we filter tables based on keywords related to the models (e.g., "gemini," "gpt"), which significantly reduces irrelevant data and minimizes LLM API usage. Moreover, we focus on leaderboard tables, which present the main results on specific benchmarks (Kardas et al., 2020), to assess LLM performance. We employ an open-source LLM (Dubey et al., 2024) to filter out non-leaderboard tables from
@@ Appendix B Details in Human Evaluation
When scoring, annotators used references from the web or literature searches to ensure the evaluation was well-informed.
Appendix C Details in Meta Analysis
@@ C.1 Which Categories Benefit from CoT?
Extracting records for this type of analysis is challenging due to the strict criteria for sample selection. While Sprague et al. (2024)’s dataset contains a larger overall sample size, as it is not limited to a single model and our dataset is not specifically oriented to replicating their study. For the overlapping model (GPT-4), their dataset includes 168 instances, whereas ours contains 553. This demonstrates the scalability of our pipeline.
@@
Category Mean $\Delta$ One-Sided Bootstrap
p-value Significant
Math 14.61 0.0000 Yes
Symbolic and algorithmic 8.85 0.0002 Yes
Spatial and temporal reasoning 3.03 0.0166 No
Logical reasoning 2.39 0.0084 No
Commonsense reasoning 5.41 0.0450 No
Multi-hop QA 2.05 0.0000 Yes
Context-aware QA 2.45 0.0014 Yes
Encyclopedic knowledge 2.18 0.0076 No
Generation 4.24 0.0280 No
Text classification 1.01 0.4464 No
Entailment 0.81 0.2070 No
Table 7: Statistical test results across different categories for replicating the study from Sprague et al. (2024). Mean $\Delta$ represents the average of improvement when using CoT over standard prompting, and significance is determined at $p=0.00227$ applying a Bonferroni correction.
@@
For the statistical significance test of each category whether the performance improved (i.e. mean improvement > 0), we perform a one-sided bootstrapping test following  Sprague et al. (2024). The results indicate significant improvements in mathematical and symbolic reasoning tasks, with a p-value of $0.00227$ with Bonferroni correction. The complete table of results is provided in Table 7.
@@ Appendix D Dataset Examples
We provide a curated selection of dataset examples in Table 8 and Table 9.
@@
Table 8: Curated list of datasets (Target Model: GPT-4 (Achiam et al., 2023)).
@@
Table 9: Curated list of datasets (Target Models: GPT-4o (Hurst et al., 2024), Claude3 Opus (Anthropic, 2024), and Gemini1.0 Pro (Team et al., 2023)).

### v2

@@ 4.2 Which Tasks Benefit from ICL?
Previous research has systematically explored ICL behavior (Min et al., 2022; Agarwal et al., 2024; Wei et al., 2023; Bertsch et al., 2024). In line with our study in section 4.1, we analyze the improvement of using in-context examples compared to not using any. To achieve this, we extract instances from LlmEvalDB that come from the same papers and tables comparing few-shot and zero-shot setups under identical conditions (i.e., model, dataset, subset, metric, and prompting method). No manual labor was needed beyond implementation, as the number of in-context examples is an integer value that can be easily computed for filtering. We measure performance deltas between few-shot and zero-shot setups, where positive values indicate few-shot advantages and negative values indicate the reverse.
Figure 7: Few-shot learning significantly improves performance over zero-shot learning in coding and multimodal tasks, whereas its impact on math tasks is less pronounced compared to CoT, as shown in Fig. 5. Grey dots represent individual deltas (improvements), sky pink dots show the mean delta per paper, and a pink star marks the mean delta for each category.
@@
Figure 7 presents the results. In contrast to CoT’s strong performance in mathematical reasoning, ICL shows only modest benefits for math tasks. However, ICL demonstrates more substantial improvements in coding and multimodal applications despite considerable performance variability across different cases. This variance may be due to the broad scope of multimodal tasks, including speech and image processing. ICL shows more uniform improvements over zero-shot performance across different categories compared to the study in Fig. 5 (std: 8.3). Note that we did not account for the number of in-context examples used; we only distinguished between the use of in-context examples and zero-shot examples. We also plot the performance improvement distribution of more versus fewer in-context examples in appendix C and find a similar overall median. This suggests that the presence of demonstrations is more important than their quantity.
@@ 4.3 Which Dataset Characteristics Negatively Affect CoT and ICL?
Motivation and Setup
To comprehend the patterns of performance degradation, we analyze cases where performance declined: specifically, where CoT resulted in worse outcomes than direct prompting and where few-shot learning performed worse than zero-shot learning (i.e. deltas below zero). Table 4 summarizes the key dataset characteristics associated with performance declines under both approaches based on analyzing the descriptions.
@@
Out of the cases of performance decline, tasks requiring expert-level knowledge showed the highest ratio for both cases, approximately 31%, indicating that knowledge-intensive tasks do not benefit much from different prompting configurations alone. Sprague et al. (2024) notes a similar point that apparent performance gains in knowledge tasks mainly stem from "reasoning" or "math" components in datasets like MMLU (Hendrycks et al., 2020). This also aligns partly with Liu et al. (2024), who found that CoT can hinder LLM performance on tasks requiring minimal deliberative thinking.
@@
Category Total Results Filtered Results
Mean $\Delta$ p-value Significant Mean $\Delta$ p-value Significant
Math 14.61 0.0000 Yes 13.53 0.0000 Yes
Symbolic and algorithmic 8.85 0.0002 Yes 9.13 0.0000 Yes
Spatial and temporal reasoning 3.03 0.0166 No 2.07 0.0056 No
Logical reasoning 2.39 0.0084 No 1.18 0.3776 No
Commonsense reasoning 5.41 0.0450 No 5.61 0.0748 No
Multi-hop QA 2.05 0.0000 Yes 1.21 0.0064 No
Context-aware QA 2.45 0.0014 Yes 0.28 0.7060 No
Encyclopedic knowledge 2.18 0.0076 No 3.31 0.0138 No
Generation 4.24 0.0280 No 1.92 0.3920 No
Text classification 1.01 0.4464 No 0.27 0.8644 No
Entailment 0.81 0.2070 No 0.93 0.1666 No
Table 7: Statistical test results across different categories for replicating the study from Sprague et al. (2024). Total Results refers to the results without filtering any papers. Filtered Results refer to the results from the filtered papers that are published in peer-reviewed journals & conferences. Mean $\Delta$ shows the average improvement when using CoT over standard prompting, and significance is determined at $p=0.00227$ after applying a Bonferroni correction.
4.4 How do CoT and ICL Interact?
Motivation and Setup
Building on previous studies section 4.1 and section 4.2, our analysis uncovers how CoT reasoning interacts with demonstrations in prompting, highlighting their complementarity. We extract instances from LlmEvalDB where CoT is used with varying levels of demonstrations, as well as instances that compare CoT and direct prompting under different demonstration levels (zero-shot vs. few-shot), all while keeping other attributes constant.
Analysis
From Table 5, demonstrations consistently enhance CoT performance: few-shot CoT outperforms zero-shot CoT by 3.0 points, and increasing the number of demonstrations yields a similar 3.1-point improvement. These comparable gains suggest that the mere presence of demonstrations, rather than their quantity, drives the improvement. However, from Table 6, we can see that when comparing CoT against standard prompting, the benefits remain relatively constant regardless of demonstration use. CoT’s improvement over standard prompting is similar in both few-shot and zero-shot settings. This indicates that while demonstrations enhance both approaches, they do not amplify CoT’s relative advantage over standard prompting.
4.5 How Do Trends in Peer-Reviewed Papers Compare to Those in arXiv Publications?
Motivation and Setup
LlmEvalDB includes information from papers published on arXiv. To examine whether the trends observed in our general analysis align with those in more selective venues, we conducted a focused analysis on a subset of peer-reviewed papers accepted at journals or conferences, using metadata from DBLP. [^3] https://dblp.org/ We report statistical significant test results from 4.1 using a subset of peer-reviewed papers. Following Sprague et al. (2024), we perform a one-sided bootstrap test to assess whether each category shows performance gains (mean improvement > 0).
Analysis
Table 7 represents the results. The statistically significant test results from our Total Results and Filtered Results indicate that math and symbolic reasoning tasks derive substantial benefit from using Chain-of-Thought (CoT) prompting, which aligns with both manual analysis (Sprague et al., 2024) and our examination in section 4.1. We can see that the core finding remains consistent even if we used a subset of peer-reviewed papers. This invariance in analysis patterns extends to other examinations like those in section 4.4, with additional details provided in Appendix C. These consistent results across both the complete dataset and the filtered selection of high-quality published papers serve to corroborate our findings and implicitly confirm our hypothesis.
@@ 5 Related Work
Previous works have focused on extracting basic result tuples (e.g., task, dataset, metric) from scientific literature (Singh et al., 2019; Hou et al., 2019; Kardas et al., 2020; Yang et al., 2022; Bai et al., 2023; Singh et al., 2024; Şahinuç et al., 2024; Kabongo et al., 2024). Our extraction pipeline improves upon this approach in two significant ways: it extracts enriched tuples that include prompting-related attributes and generates detailed dataset descriptions by leveraging LLM and automatically linked source papers. Hence, unlike previous works that primarily compiled leaderboard tables, our enhanced extraction pipeline enables deeper review analysis, contributing to the broader goal of AI-driven scientific discovery (Xu et al., 2021; Majumder et al., 2024; M. Bran et al., 2024).
@@
Our study focuses on extracting experimental results of frontier proprietary LLMs (Achiam et al., 2023; Anthropic, 2024; Team et al., 2023), with a specific emphasis on target attributes that incorporate information about prompting methods (Brown et al., 2020; Wei et al., 2022; Min et al., 2022). In the context of prompting, prior studies have systematically analyzed the mechanisms behind prompting methods, focusing either on the use of in-context examples (Min et al., 2022; Lampinen et al., 2022; Weber et al., 2023; Zhang et al., 2022) or techniques that elicit reasoning, such as CoT prompting (Wei et al., 2022; Shaikh et al., 2023; Wang et al., 2023; Turpin et al., 2024; Sprague et al., 2024; Liu et al., 2024). Conversely, we examine the model’s behavior by conducting a literature analysis, which compiles data from scientific sources to reveal insights.
Literature Analysis
Literature analysis systematically aggregates and examines data from multiple independent studies on a given topic to derive more precise and reliable conclusions. It has been widely applied in the biomedical domain for identifying target materials or clinical records (Bao et al., 2019; Yun et al., 2024). In the NLP domain, review analysis has been used for metric standardization (Reiter, 2018), literature review (Santu et al., 2024; Du et al., 2024), and assessing evaluation criteria for specific domains (Ostheimer et al., 2023). In contrast, our work employs a review analysis approach to evaluate the behavior of LLMs. In the context of LLMs, Asai et al. (2024) utilizes retrieval-augmented language models to synthesize scientific literature. However, this approach can only process a limited number of documents during retrieval to synthesize. The work by Sprague et al. (2024) is perhaps the most closely related to ours. They conducted a review analysis through a literature survey to examine the effectiveness of CoT prompting for LLMs. However, their study is focused on CoT prompting, conducted on a limited scale, and relies on manual extraction methods.
@@ 6 Conclusion
Our study streamlines literature analysis by using an LLM for dataset extraction, demonstrating that the automatically extracted dataset, LlmEvalDB, can yield novel findings and replicate manual analyses. We confirm the dataset’s quality by replicating a key finding from Sprague et al. (2024). Our analysis provides insights into prompting configurations, such as the benefits of ICL, the combined effects of CoT and ICL, and the dataset characteristics on performance declines with CoT or ICL. Overall, our resources support ongoing literature analyses, enhancing the understanding of LLM behaviors.
@@ 7 Limitations
Our study focuses on four leading proprietary LLMs selected for their widespread adoption and extensive documentation in existing literature. We excluded newer models like GPT4-o1 and Deepseek-R1 due to limited published results, though our analysis pipeline can easily accommodate them as more experimental data becomes available. While most analyzed models are not fine-tunable, this limitation presents an opportunity for future research combining fine-tuning and prompting methods analysis.
@@
Our literature analysis aims to identify trends by aggregating information across multiple studies, which generates potential hypotheses but requires systematic validation. While we report findings and patterns observed in the aggregated scientific literature, we did not independently validate each claim or finding. This limitation suggests the need for future work to rigorously test the hypotheses emerging from our analysis.
@@ Acknowledgments
We would like to thank Microsoft’s Azure Accelerate Foundation Models Research Program and NVIDIA’s Academic Grant Program for providing computational resources to support this work. This research is supported in part by the NSF under grant numbers IIS-2052498 and SMA-2418946. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. We also appreciate Ethan Mendes, Duong Minh Le, and Seongeun Park for their valuable discussions and feedback on the paper.
@@ Appendix A Details and Trials During Extraction
Hyperparameters and Prompt Tuning
We did not extensively tune any LLM hyperparameters, instead, we used OpenAI’s default generation setup with greedy decoding. For prompting, we initially built upon previous work (Bai et al., 2023) as a foundation. Still, we adapted the prompts to better suit our project’s objectives, such as (1) determining whether the table includes experimental results for our target model or (2) enhancing records using contextual information from the paper.
@@
Before the extraction stage, we filter tables to reduce extraction time and API costs. To cut computational expenses, we pre-filter tables for those containing results of target models. Using simple heuristics, we filter tables based on keywords related to the models (e.g., "gemini," "gpt"), which significantly reduces irrelevant data and minimizes LLM API usage. Moreover, we focus on leaderboard tables, which present the main results on specific benchmarks (Kardas et al., 2020), to assess LLM performance. We employ an LLAMA3.1-70B-Instruct (Dubey et al., 2024) to filter out non-leaderboard tables.
@@ Appendix B Details in Human Evaluation
When scoring, annotators used references from the web or literature searches to ensure the evaluation was well-informed. Our scoring rubric is designed to approximate an interval scale, allowing us to compute average scores. This approach aligns with standard practices in the machine learning field for evaluating response quality, whether through model-based or human-based assessments (Liu et al., 2023b; Kim et al., 2023).
Appendix C Details in Literature Analysis of Prompting Behavior in Frontier LLMs
@@ C.1 Which Categories Benefit from CoT?
Extracting records for this type of analysis is challenging due to the strict criteria for sample selection. While Sprague et al. (2024)’s dataset contains a larger overall sample size, as it is not limited to a single model and LlmEvalDB is not specifically oriented to replicating their study. For the overlapping model (GPT-4), their dataset includes 168 instances, whereas ours contains 553. This demonstrates the scalability of our pipeline.
@@
CoT Reasoning Original Results Filtered Results
Median Q1 Q3 Median Q1 Q3
$\Delta$ (Few-shot, Zero-shot) 3.0 0.4 9.2 5.0 2.3 10.4
$\Delta$ (More-shot, Less-shot) 3.1 0.4 9.0 5.0 2.3 10.7
Table 8: Statistical distribution of performance improvements when using varying levels of demonstrations in CoT prompting. Original Results refers to the results from Table 5 from the submission (without filtering any papers). Filtered Results refer to the results from the filtered papers that are published in top conferences. $\Delta$(A, B) refers to the performance improvement of A over B, Q1 and Q3 represent the first and third quartiles, respectively.
$\Delta$ (CoT, Direct Prompting) Original Results Filtered Results
Median Q1 Q3 Median Q1 Q3
Few-shot 0.9 -1.2 3.7 1.3 0.0 3.0
Zero-shot 1.3 -0.4 4.7 1.3 0.0 3.6
Table 9: Statistical distribution of performance improvements for CoT compared to direct prompting given different levels of demonstrations. Original Results refers to the results from Table 6 from the submission (without filtering any papers). Filtered Results refer to the results from the filtered papers that are published in top conferences.
@@ C.2 Which Categories Benefit from ICL?
C.3 How Do Trends in Peer-Reviewed Papers Compare to Those in arXiv Publications?
We reanalyzed the joint behavior findings from section 4.4 using only peer-reviewed papers published in journals or conferences. Table 8 and Table 9 present these results. The filtered data confirms our original finding regarding the interaction between CoT and ICL: CoT with demonstrations (ICL) consistently outperforms CoT without demonstrations. Additionally, CoT’s improvement over standard prompting remains consistent regardless of demonstration count (zero-shot or few-shot).
@@ Appendix D Dataset Examples
We provide a curated selection of examples from LlmEvalDB in Table 10 and Table 11.
@@
Table 10: Curated list of examples from LlmEvalDB (Target Model: GPT-4 (Achiam et al., 2023)).
@@
Table 11: Curated list of examples from LlmEvalDB (Target Models: GPT-4o (Hurst et al., 2024), Claude3 Opus (Anthropic, 2024), and Gemini1.0 Pro (Team et al., 2023)).